the product of tanay's age five yrs ago and his age ten yrs later is 16.find their present agd.
Answer by prateekagrawal(56)

Let Tanny's present age be x years..
so, his age before five years = (x-5) years
and his age after ten years = (x+10) years
hence,
according to given condition ,
(x-5)(x+10)=16
x^2-5x+10x-50=16
x^2+5x-50-16=0
x^2+5x-66=0
x^2+11x-6x-66=0
x(x+11)-6(x+11)=0
(x+11)(x-6)=0
x=-11 or x=6
since age can not be negative , so x=6
hence , Tanny's age = 6 years . :)